Please help me to solve this equation.
2x + 1/2+ = x - 3/4

Answer by malakumar_kos@yahoo.com(315)



solve this equation.
2x + 1/2+ = x - 3/4
2x+1/2 = x-3/4

2x+x = -3/4-1/2
3x = -3-2/4
3x = -5/4
x = -5/4.3 = -5/12